acopeland27:

Maria has a savings account that is earning simple interest. If she started with $200 in the account and has earned a total of $8 in interest in two years, what is the rate of the account? Use the formula .

sammixboo:

We are going to use the interest formula again for this question too! SI = P x R x T Our principle is $200, total is $8, and time is 2. Let's plus this into the equation and solve: 8 = 200(2r) Do you know how to solve for r to get our answer?

acopeland27:

no

acopeland27:

im only a seventh grader

sammixboo:

That's okay! So we solve this equation by using PEMDAS but in reverse, to isolate the r onto one side of the equation. To start with this, we need to divide 200 on both sides 8/200 = 200(2r)/200 What is 8/200?

sammixboo:

And oh interesting, I am surprised you are learning this in 7th grade!

acopeland27:

0.04

acopeland27:

did he leave

sammixboo:

Yep! So now we have 0.04 = 2r Do you know what we do next to solve for r?

acopeland27:

no

acopeland27:

im sorry

acopeland27:

:(

sammixboo:

You're okay! We have to divide 2 on both sides of the equation. So 0.04/2 = 2r/2 What is 0.04/2? that will be our answer. Also, were you ever taught PEMDAS? it seems weird that you're learning interest so early!

acopeland27:

yes but doing it with terms are hard

acopeland27:

0.02

sammixboo:

Correct! So the interest rate is 0.02 or 2%

acopeland27:

oh thankyou.